OpenStudy (anonymous):

State the slope and the y-intercept for the graph of the equation. 3x + 6y = -6

sam (.sam.):

If you can arrange it in this form: y=mx+b, m is the slope b is the y-intercept

OpenStudy (anonymous):

don't I have to get the y by itself?

sam (.sam.):

Yes you have to

OpenStudy (boldgeorge):

\[y = \frac{ 1 }{ 2 } (-x-2)\]

OpenStudy (anonymous):

wait whats the work for it? Do I subtract it like this: 3x + 6y = -6 =3x =3x

OpenStudy (anonymous):

m=-1/2 b=-1 m=slope b=y-int

sam (.sam.):

\[3x + 6y = -6\] Subtract 3x both sides \[6y=-3x-6\] Divide 6 both sides \[y=-\frac{1}{2}x-1\] ------------------------ m is -1/2 b is -1

OpenStudy (anonymous):

thank you :) how do I graph it?

OpenStudy (anonymous):

this?

OpenStudy (anonymous):

?

sam (.sam.):

Yes

OpenStudy (anonymous):

how i did it was ok?

OpenStudy (anonymous):

yay, thanks :) can you help me with the next problem?

sam (.sam.):

If your plot includes negative values of x that'd be greater

OpenStudy (anonymous):

Write an equation in slope-intercept form to represent the table values. 

OpenStudy (anonymous):

@.Sam.

sam (.sam.):

Equation construction \(y-y_1=m(x-x_1)\) We need a coordinate \((x_1,y_1)\) and slope, m. --------------------------------------- To find slope Take the coordinates \((x_1,y_1)\) and \((x_2,y_2)\) Slope formula is \[Slope=\frac{y_2-y_1}{x_2-x_1}\] Substitute and find slope ---------------------------------------

OpenStudy (anonymous):

can I choose any coordinates?

OpenStudy (anonymous):

I got 7